OpenStudy (anonymous):

Find the ratio in simplest form. 30:6 1:5 4:1 5:1

OpenStudy (anonymous):

i think its 6:1

OpenStudy (mayankdevnani):

30/6=5/1=5:1

OpenStudy (mayankdevnani):

ok @lawson4238

OpenStudy (anonymous):